Lebron James plays for the Cleveland Cavaliers and shoots on average twenty one times per game. Out of those twenty one shots he makes nine on average. If Lebron took three hundred and fifty seven shots for the month of November how many shots did he miss? What percent of the total shots were misses? Makes?
Answer by Theo(2971)

average game he makes 9 out of 21.

His make percentage is 9/21 = .428571429 * 100% = 42.8571429%

He took 357 shots for the month of November.

Assuming he hits the average in all the games, then he hits .428571429 * 357 = 153 and he misses 357 - 153 = 204
This means he misses 204/357 = .571428571 * 100% = 57.1428571% of the time.